Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
bersanetti, ruggi, tapia - 17:59 Monday 01 March 2021 (50929) Print this report
CITF Acl process preparations and CITF alignments

This morning we started trying to solve issues related with DAQ together with Alain, as reported in logbook entry #50920. The Tolm Packet size was increased and then we could send to the DAQ all the channels we needed.

We used this time to add and implement a new configuration file for SRCL (/virgoData/VirgoOnline/LSC_SRCL.cfg) in the LSC_Acl process. As of now, the file contains the usual loop structure, recycled from PRCL, with the needed changes for the sensing matrix and possible trigger signals that SRCL will use for its single lock acquisition, as it is done for PRCL. This is for the single recycling cavity lock and auxiliary commissioning configurations. The integration with the new trigger logic will be done soon.

Then, together with Paolo, we started with the CITF alignment in a safely way. With the arms locked on IR, SR parked and PR parked. We first misaligned NE, WE and WI. B1p was very quiet, so we scan few urad for PR and NI to see a clear movement on B1p camera. Next, leaving PR parked and NI in its good position, we started moving SR to get power built between NI and SR. Finally we fine adjust SR in Yaw and Pitch to have multiple reflections centered (GPS: 1298629178 + 300 sec)

  •  UTC 10:19:20 PR misalinged and scans of SR. See figure 1. CITF_SR_scans_20210301.png

While doing all this, we implemented the new code for the Metatron Nodes for the Suspensions, in order to use and also detect the new safety. The tests went fine, here some more remarks:

  •  MAR_[TX,TY]_SET setpoint MUST NOT be used for the parking of the mirrors; instead, the new gNAME MAR_TY_ALI should be used for this; this is what Metatron does now, the "delta" feature has been decommissioned.
  • All SUS nodes support as usual the SAVE_ALIGNED feature, but please do not use it for the PR just yet, we have to understand a bug, possibly related to a too fast read/write sequence which happens only in the PR case; given the new structure, SAVE_MISALIGNED will do nothing for PR and SR.
  • UTC 10:28:05  SR MAR TX was misaligned in order to trigger PR to move to a safe position.  See figure 2. CITF_PR_security_test_20210301.png
  • SUS_PR was able to detect the safety tripping; the procedure is the following: it detects the mirror being parked and it self-requests the MISALIGNED state, forcing the new gNAME accordingly; from there onwards the suspension is simply misaligned, but with a notification message explaining the tripping of the safety; one can realign the mirror, but in this case it is best to check things by hand. The same will be replicated for the SR mirror, but this is not online yet.
  • UTC 10:48:30 start misaligning SR and PR, then lock again arms on IR. See figure 3. CITF_SR_PR_parked_arms_locked_20210301.png

At around 11:30 UTC we waited for DAQ/Computing to recover from the crash of rtpc7, and the check of the Injection.

At 12:10 UTC we locked the arms again on IR, but we could see that the power in transmission in B7 and B8 was decreasing over time. Analizing the FFTs of the transmitted power, we could see peaks of noise at around 7 Hz and 9 Hz (one per arm), which are close to the dithering lines frequencies for NArm and WArm. After noticing this we unlocked the arms. We cheked if the drift control was engaged, when the comand was sent using the corresponding Metatron node, but it was not the case, but only for the Acl side. Therefore we rebooted the ASC_Acl process and we tried again to turn the boost on for NArm and WArm, and we recovered the cavities.


Moving forward, at 14:20 UTC we started the CITF alignment in order to collect more data for the search of triggers. For clarity, here we describe the steps of the alignment.

  • WE, WI and NI misaligned using NARM and WARM metatron nodes.
  •  PR aligned from Tx = -43 urad Ty = -49.9 urad to Tx = -43 urad Ty = 0.4 urad and then was fine-tuned to Tx = -38.5 urad Ty = -0.7 urad looking at B1p camera. Then we saved these values using metatron's SAVE_ALIGNED state of SUS_PR.
  •  PR misaligned to Tx = -38.5 urad Ty = -51 urad using PR metatron node (MISALIGNED state).
  •  SR aligned from Tx = -84 urad Ty = -671.5 urad to Tx = 84 urad Ty = -193.5 urad and then was fine-tuned to Tx = 83.9 urad Ty = -192.3 urad looking at B1p camera. Then we saved these values using metatron's SAVE_ALIGNED state of SUS_SR.
  • PR realigned using PR metatron node (ALIGNED state), which uses previously saved values in step 2.
  • WI aligned from Tx = 67 urad and Ty = 100 urad to Tx = 13.6 urad and Ty = -8.5 rad.

We left the Interferometer alone for 10 min at UTC 15:04. All this process can be seen in figure 4.

At the end of the shift we restarted LSC_Acl, after implementing many changes:

  • we acquired some new channels, all from SIB2_Photodiodes; as they were already present in the Tolm packet, there was no issue about it;
  • we added the demodulation phase monitor for B2_56MHz;
  • we generated a few normalized signals for sensing and triggers for the CITF, and sent them to the DAQ;
  • we integrated the new LSC_SRCL configuration file, and added a first draft for the lock flags, and sent all the new relevant channels to the DAQ;

Please remember that as of now there is no automation about SRCL, so any change done by hand has to be reverted by hand as well.

Images attached to this report
Search Help